OpenStudy (anonymous):

Yuri solved the equation below incorrectly. Explain the step where he went wrong. fifteen times b plus one equals thirty-one Step 1: fifteen times b plus one minus one equals thirty-one minus one Step 2: fifteen times b equals thirty Step 3: fifteen times b times fifteen equals thirty times fifteen Step 4: b equals four hundred fifty

OpenStudy (anonymous):

i think the first step

OpenStudy (mertsj):

\[15b+1=31\] \[15b+1-1=31-1\] \[15b=30\] \[\frac{15b}{15}=\frac{30}{15}\] \[b=2\]

OpenStudy (anonymous):

thanks

OpenStudy (mertsj):

yw